What is the Domain of the function y = (x 2 – 2) /( x – 2).
Also, find f(x + 1).

Answer by stanbon(75887)

What is the Domain of the function y = (x^2 – 2) /(x – 2).
Domain: All Real Numbers except x = 2
------------------------------------------
Also, find f(x + 1).
f(x+1) = [(x+1)^2-2)/(x+1-2)]
f(x+1) = [x^2 + 2x -1]/[x-1]
========================
Cheers,
Stan H.